# ChaosBench: A Multi-Channel, Physics-Based Benchmark for Subseasonal-to-Seasonal Climate Prediction

### v3

@@ Abstract
Accurate prediction of climate in the subseasonal-to-seasonal scale is crucial for disaster preparedness and robust decision making amidst climate change. Yet, forecasting beyond the weather timescale is challenging because it deals with problems other than initial conditions, including boundary interaction, butterfly effect, and our inherent lack of physical understanding. At present, existing benchmarks tend to have shorter forecasting range of up-to 15 days, do not include a wide range of operational baselines, and lack physics-based constraints for explainability. Thus, we propose ChaosBench, a challenging benchmark to extend the predictability range of data-driven weather emulators to S2S timescale. First, ChaosBench is comprised of variables beyond the typical surface-atmospheric ERA5 to also include ocean, ice, and land reanalysis products that span over 45 years to allow for full Earth system emulation that respects boundary conditions. We also propose physics-based, in addition to deterministic and probabilistic metrics, to ensure a physically-consistent ensemble that accounts for butterfly effect. Furthermore, we evaluate on a diverse set of physics-based forecasts from four national weather agencies as baselines to our data-driven counterpart such as ClimaX, PanguWeather, GraphCast, and FourCastNetV2. Overall, we find methods originally developed for weather-scale applications fail on S2S task: their performance simply collapse to an unskilled climatology. Nonetheless, we outline and demonstrate several strategies that can potentially extend the predictability range of existing weather emulators, including the use of ensembles and robust control of error propagation. Our benchmark, datasets, and instructions are available at https://leap-stc.github.io/ChaosBench.
@@ 1 Introduction
Although critical for economic planning, disaster preparedness, and policy-making, subseasonal-to-seasonal (S2S) prediction is lagging behind the more established field of short/medium-range weather, or long-range climate predictions. For instance, many natural hazards tend to manifest in the S2S scale, including the slow-onset of droughts that lead to wildfire [1, 2], heavy precipitations that lead to flooding [3], and persistent weather anomalies that lead to extremes [4]. So far, current approaches to weather and climate prediction are heavily reliant on physics-based models in the form of Numerical Weather Prediction (NWP). Many NWPs are based on the discretization of governing equations that describe thermodynamics, fluid flows, etc. However, these models are expensive to run especially in high-resolution setting. For example, there are massive computational overheads to perform numerical integration at fine spatiotemporal resolutions that are operationally-useful [5]. Furthermore, their relative inaccessibility to non-experts is a major roadblock to the broader community. As a result, there is a growing interest to apply data-driven models to emulate NWPs, as they tend to have faster inference speed, are less resource-hungry, and more accessible [6, 7, 8, 9, 10, 11, 12]. Nevertheless, many data-driven benchmarks have so far been focused on the short (1-5 days), medium (5-15 days), and long (years-decades) forecasting ranges. In this work, we include S2S as a more challenging task that requires different emulation strategies: being in between two extremes, it is doubly sensitive to (1) initial conditions (IC) as in the case for short/medium-range weather, and (2) boundary conditions (BC) as in the case for long-range climate [13, 14, 15].
@@
We propose ChaosBench to bridge these gaps (Figure 1). It is comprised of variables beyond the typical surface-atmospheric ERA5 to also include ocean, ice, and land reanlysis products that span over 45 years to allow for full Earth system emulation that respects boundary processes. We also provide 44-day ahead physics-based control (deterministic) and perturbed (ensemble) forecasts as baselines from four national weather agencies over the last 8 years. In addition, we introduce physics-based and incorporate probabilistic-based, in addition to deterministic metrics, for a more physically-consistent ensemble that accounts for butterfly effect. As far as we know, ChaosBench is one of the first to systematically evaluate several state-of-the-art data-driven models including ViT/ClimaX [16], PanguWeather [17], GraphCast [7], and FourCastNetV2 [9] on S2S predictability.
@@
Figure 1: We propose ChaosBench, a large-scale, fully-coupled, physics-based benchmark for subseasonal-to-seasonal (S2S) climate prediction. It is framed as a high-dimensional sequential regression task that consists of 45-year, multi-system observations for validating physics-based and data-driven models, and training the latter. Physics-based forecasts are generated from four national weather agencies with 44-day lead-time and serve as baselines to data-driven forecasts. Our benchmark is one of the first to incorporate physics-based metrics to ensure physically-consistent and explainable models. The blurred image at $\Delta t=44$ represents a challenge of long-term forecasting.
@@ 2 Related Work
In recent years, several benchmarks have been introduced to push the field of data-driven weather and climate prediction [18, 19, 20, 21, 22, 23, 24]. We analyze the limitations of existing works, and propose how ChaosBench fills in these gaps (see Table 1, Appendix C).
@@
Limited spatiotemporal extent. Many S2S benchmarks tend to focus on regional forecasts, such as the US [18, 25]. In addition, the temporal extent of observation with common interval is more varied, with some less than 20 years [19, 23]. ChaosBench has the most extensive overlapping temporal coverage yet, extending to 45 full years of inputs covering multiple reanalysis products beyond ERA5.
@@
Lack of physics-based constraints. So far, limited number of benchmarks have explicitly incorporated physical principles to improve or constrain forecasts. ChaosBench introduces physics-based metrics that can be easily used for comparison (scalar) and integrated into the training/inference pipeline (differentiable).
@@
SubseasonalRodeo [18] 11 2 44 western US
SubseasonalClimateUSA [25] 25 2 44 contiguous US
CliMetLab [23] 20 2 44 global
@@
Table 1: Comparison with other benchmark datasets: ChaosBench (ours) is evaluated on the largest set of global variables, benchmarked against the largest number of operational NWPs (four national agencies in the US, Europe, UK, and Asia), and incorporates both physics-based and probabilistic metrics for a more physically-consistent S2S ensemble forecast.
@@ 3.1 Observations
We discuss the components of ChaosBench, including the global reanalysis products of surface-atmosphere (ERA5), terrestrial (LRA5), and sea-ice (ORAS5), as well as simulations from physics-based models. The spatiotemporal resolutions of the former are matched with the latter’s daily forecasts at $1.5^{\circ}$ to allow for consistent evaluation and integration e.g., hybrid physics-based emulator.
ERA5 reanalysis provides a comprehensive record of the global atmosphere combining physics and observations for correction [26]. We processed their hourly data from 1979 to 2023 and selected measurements at the 00UTC step. The variables include temperature ($t$), specific humidity ($q$), geopotential height ($z$), and 3D wind speed ($u,v,w$) at 10 pressure levels: $1000,925,850,700,500,300,200,100,50,10$ hpa, totalling 60 variables.
LRA5 or ERA5-Land reanlysis provides a detailed record of variables governing global terrestrial processes with specific corrections tailored for land surface applications such as flood forecasting [27] or carbon fluxes [21, 28]. We processed hourly data from 1979 to 2023 and selected measurements at the 00UTC step, for a total of 43 variables, including t2m, u10, v10, and tp (full list in D.1.1).
ORAS5 or the Ocean Reanalysis System 5 provides an extensive record of sea-ice variables that incorporate multiple depth levels [29]. Since the public data is available on a monthly basis, we replicate them for daily compatibility with temporal extent between 1979 and 2023, for a total of 21 variables, including sst and ssh (full list in D.1.2).
@@ 3.3 Auxiliary
Figure 3: Motivating problem: as we perform longer rollouts, the (a) residual error becomes larger and the prediction becomes blurry. This behavior is captured in the Fourier frequency domain where the (b) power spectra $S(k)$ at low wavenumber $k$ (i.e., low frequency signal) remains consistent at long rollouts, but not for higher $k$ (i.e., high frequency signal). This phenomenon explains why long-term forecasts excel at capturing large-scale pattern but not fine-grained details i.e., smooth.
@@ 4.2 Physics Metrics
As illustrated in Figure 3, we find that in general, data-driven forecasts tend to become blurry (Figure 3(a)) due to power decay in the spectral domain 3(b). This motivates us to propose two physics-based metrics that measure the deviation or difference between the power spectra of prediction ${\hat{S}}(k)$ and target $S(k)$, where $k\in\mathbf{K}$, and $\mathbf{K}$ is the set of all scalar wavenumbers from 2D Fourier transform. Focusing on high-frequency components, we introduce $\mathbf{K}_{q}=\{k\in\mathbf{K}\mid k\geq Q(q)\}$, where $Q$ is the quantile function of $\mathbf{K}$ and $q\in[0,1]$. We set $q=0$ or $q=0.9$ for training and evaluation respectively. We denote $S_{q}=\{S(k)\mid k\in\mathbf{K}_{q}\}$ as the corresponding power spectra on $\mathbf{K}_{q}$, and we normalize the distribution to $S^{\prime}(k)$ such that it sums up to 1. Similarly we use $\hat{S}^{\prime}(k)$ to denote the normalized power for predictions.
@@ 4.3 Probabilistic Metrics
(a) RMSE ($\downarrow$ is better)
(b) MS-SSIM ($\uparrow$ is better)
Figure 4: Evaluation results between baseline climatology (black lines) and physics-based control/deterministic forecasts. At longer forecasting horizon, most physics-based control forecasts perform worse than climatology. This suggests knowledge gap, such as poorly represented subgrid processes.
(a) RMSE ($\downarrow$ is better)
(b) MS-SSIM ($\uparrow$ is better)
Figure 5: Evaluation results between baseline climatology (black lines) and data-driven models including PanguWeather (PW), GraphCast (GC), and FourCastNetV2 (FCN2). We observe that data-driven models perform worse than climatology on S2S timescale. Note: FCN2 lacks q-700.
5 Benchmark Results
Throughout this section, we report headline results on $\hat{\mathbf{X}}\in\{T850,Z500,Q700\}$, following Weatherbench 2 [37]. The full benchmark scores are available at https://leap-stc.github.io/ChaosBench. We primarily use four state-of-the-art models for comparison including ViT/ClimaX [16], PanguWeather [17], GraphCast, and FourCastNetV2 [9] [7]. However, whenever ablation is performed, we use popular baselines including Lagged Autoencoder [38], ResNet [39], UNet [22], and FNO [40] trained on 1979-2015 data and validated on 2016-2021 data. All evaluations presented here are done on the held-out 2022-2023 data. The full implementation details are discussed in Appendix E.
Ensemble Forecasts Account for Initial Condition Uncertainty. Despite the underperformance of deterministic models, many studies have highlighted the potential of ensemble forecasts to account for trajectory divergences caused by IC uncertainties, also known as the butterfly effect [15]. Figure S4 shows that the performance of ensembles across physics-based models improves relative to their deterministic counterparts. For instance, when we take the metrics ratio between ensemble and deterministic forecasts as in Figure 7 (+ S5), the ratio of RMSE decreases with lead time, while the ratio of MS-SSIM improves over time with little significant changes in SpecDiv. This highlights the importance of building a well-dispersed ensemble that accounts for long-range divergences for improved S2S predictability.
@@ 5 Benchmark Results
Figure 7: Metrics ratio e.g., $\texttt{RMSE}_{ens}/\texttt{RMSE}_{det}$ between ensemble and deterministic forecasts, where the former improves the latter by accounting for IC uncertainty that can lead to trajectory divergences.
Figure 8: Probabilistic evaluation on ensemble forecasts indicating current skill limits of 15-20 days; CRPSS $>0$ suggests skills better than climatology variability.
@@
Physical Constraints Yield Improved Performance. We find models that explicitly incorporate physical knowledge (e.g., learning spectral signals, beyond pixel information) have better performance across metrics, such as FNO, as summarized in Table S7 given identical parameter budget of $10^{6}$.
Current Limits of S2S Predictability. Given our best models, we evaluate the extent of predictability in order to base our next steps. As illustrated in Figure 8 (+ S6), we find that ECMWF high-resolution ensemble, dubbed as the gold standard, still has the best performance in terms of CRPSS (vs ERA5 climatology), with a predictability range of around 15-20 days ahead before its skill collapses to climatology (i.e., CRPSS $\rightarrow$ 0). However, the resurgence of data-driven models are rapidly transforming the field as they are able to efficiently distil knowledge and automatically discover emergent patterns from large high-dimensional dataset, instead of first reducing them to physical functions with limited set of variables requiring constant calibration as is traditionally done in NWPs.
@@ 6 Conclusion
We present ChaosBench, a challenging benchmark to extend the predictability range of weather emulators into the S2S timescale where many processes with significant socioeconomic repercussions tend to occur, including extreme events. In addition to providing diverse datasets beyond ERA5 for a full Earth system emulation, we also perform extensive benchmarking on state-of-the-art data-driven and physics-based models alike. Through various ablation, we systematically find that skillfulness can potentially be extended by ensemble forecasting, controlling for exponential error growth, and incorporating physical knowledge.
Limitations. Nonetheless, our input datasets have relatively coarse spatiotemporal resolution to match that of physics-based S2S forecasts. In the future, we plan to provide multi-resolution datasets to enable for better modeling approaches.
@@
The authors would like to acknowledge funding, computing, and storage resources from the NSF Science and Technology Center (STC) Learning the Earth with Artificial intelligence and Physics (LEAP) (Award #2019625) and the Department of Energy. AG would like to acknowledge support from Google and Schmidt Sciences.
@@
Juan Nathaniel1, Yongquan Qu1, Tung Nguyen2, Sungduk Yu3,5, Julius Busecke1,4,
@@
1Columbia University, 2UCLA, 3UCI, 4LDEO, 5 Intel Labs

### v5

@@ Abstract
Accurate prediction of climate in the subseasonal-to-seasonal scale is crucial for disaster preparedness and robust decision making amidst climate change. Yet, forecasting beyond the weather timescale is challenging because it deals with problems other than initial condition, including boundary interaction, butterfly effect, and our inherent lack of physical understanding. At present, existing benchmarks tend to have shorter forecasting range of up-to 15 days, do not include a wide range of operational baselines, and lack physics-based constraints for explainability. Thus, we propose ChaosBench, a challenging benchmark to extend the predictability range of data-driven weather emulators to S2S timescale. First, ChaosBench is comprised of variables beyond the typical surface-atmospheric ERA5 to also include ocean, ice, and land reanalysis products that span over 45 years to allow for full Earth system emulation that respects boundary conditions. We also propose physics-based, in addition to deterministic and probabilistic metrics, to ensure a physically-consistent ensemble that accounts for butterfly effect. Furthermore, we evaluate on a diverse set of physics-based forecasts from four national weather agencies as baselines to our data-driven counterpart such as ViT/ClimaX, PanguWeather, GraphCast, and FourCastNetV2. Overall, we find methods originally developed for weather-scale applications fail on S2S task: their performance simply collapse to an unskilled climatology. Nonetheless, we outline and demonstrate several strategies that can extend the predictability range of existing weather emulators, including the use of ensembles, robust control of error propagation, and the use of physics-informed models. Our benchmark, datasets, and instructions are available at https://leap-stc.github.io/ChaosBench.
11footnotetext: Corresponding author: jn2808@columbia.edu
@@ 1 Introduction
Although critical for economic planning, disaster preparedness, and policy-making, subseasonal-to-seasonal (S2S) prediction is lagging behind the more established field of short/medium-range weather, or long-range climate predictions. For instance, many natural hazards tend to manifest in the S2S scale, including the slow-onset of droughts that lead to wildfire [1, 2], heavy precipitations that lead to flooding [3], and persistent weather anomalies that lead to extremes [4]. So far, current approaches to weather and climate prediction are heavily reliant on physics-based models in the form of Numerical Weather Prediction (NWP). Many NWPs are based on the discretization of governing equations that describe thermodynamics, fluid flows, etc. However, these models are expensive to run especially in high-resolution setting. For example, there are massive computational overheads to perform numerical integration at fine spatiotemporal resolutions that are operationally useful [5]. Furthermore, their relative inaccessibility to non-experts is a major roadblock to the broader community. As a result, there is a growing interest to apply data-driven models to emulate NWPs, as they tend to have faster inference speed, are less resource-hungry, and more accessible [6, 7, 8, 9, 10, 11, 12]. Nevertheless, many data-driven benchmarks have so far been focused on the short (1-5 days), medium (5-15 days), and long (years-decades) forecasting ranges. In this work, we include S2S as a more challenging task that requires different emulation strategies: being in between two extremes, it is doubly sensitive to (1) initial conditions (IC) as in the case for short/medium-range weather, and (2) boundary conditions (BC) as in the case for long-range climate [13, 14, 15, 16].
@@
We propose ChaosBench to bridge these gaps (Figure 1). It is comprised of variables beyond the typical surface-atmospheric ERA5 to also include ocean, ice, and land reanalysis products that span over 45 years to allow for full Earth system emulation that respects boundary processes. We also provide 44-day ahead physics-based control (deterministic) and perturbed (ensemble) forecasts from four national weather agencies over the last 8 years as baselines. In addition, we introduce physics-based and incorporate probabilistic, in addition to deterministic metrics, for a more physically-consistent ensemble that accounts for butterfly effect. As far as we know, ChaosBench is one of the first to systematically evaluate several state-of-the-art data-driven models including ViT/ClimaX [17], PanguWeather [18], GraphCast [7], and FourCastNetV2 [9] on S2S predictability.
@@
Figure 1: We propose ChaosBench, a large-scale, fully-coupled, physics-based benchmark for subseasonal-to-seasonal (S2S) climate prediction. It is framed as a high-dimensional sequential regression task that consists of 45+ years, multi-system observations for validating physics-based and data-driven models, and training the latter. Physics-based forecasts are generated from four national weather agencies with 44-day lead-time and serve as baselines to data-driven forecasts. Our benchmark is one of the first to incorporate physics-based metrics to ensure physically-consistent and explainable models. The blurred image at $\Delta t=44$ represents a challenge of long-term forecasting.
@@ 2 Related Work
In recent years, several benchmarks have been introduced to push the field of data-driven weather and climate prediction [19, 20, 21, 22, 23, 24, 25, 26, 27]. We analyze the limitations of existing works, and propose how ChaosBench fills in these gaps (see Table 1, more justifications in Appendix C).
@@
Limited spatiotemporal extent. Many S2S benchmarks tend to focus on regional forecasts, such as the US [23, 24]. In addition, the temporal extent of observation with common interval is more varied, with some less than 20 years [19, 25]. ChaosBench has the most extensive overlapping temporal coverage yet, extending to 45+ years of inputs covering multiple reanalysis products beyond ERA5.
@@
Lack of physics-based constraints. So far, limited number of benchmarks have explicitly incorporated physical principles to improve or constrain forecasts. ChaosBench introduces physics-based metrics that can be used for comparison (scalar) and integrated into ML pipeline (differentiable).
Table 1: Comparison with other benchmark datasets: ChaosBench (ours) is evaluated on the largest set of global variables, benchmarked against large number of operational NWPs (four national agencies in the US, Europe, UK, and Asia), and incorporates both physics-based and probabilistic metrics for a more physically-consistent S2S ensemble forecast.
@@
SubseasonalRodeo [23] $<$30 2 44 western US
SubseasonalClimateUSA [24] $<$30 2 44 contiguous US
CliMetLab [25] $<$30 2 44 global
@@ 3.1 Observations
We discuss the components of ChaosBench, including the global reanalysis products of surface-atmosphere (ERA5), sea-ice (ORAS5), and terrestrial (LRA5), as well as simulations from physics-based models. The spatiotemporal resolutions of the former are matched with the latter’s daily forecasts at $1.5^{\circ}$ to allow for consistent evaluation and integration e.g., hybrid physics-based emulator. However, we provide a one-liner script to process higher e.g., $0.25^{\circ}$ resolution input in Section B.4.
ERA5 Reanalysis provides a comprehensive record of the global atmosphere combining physics and observations for correction [28]. We processed their hourly data from 1979 to present and selected measurements at the 00UTC step. The variables include temperature ($t$), specific humidity ($q$), geopotential height ($z$), and 3D wind speed ($u,v,w$) at 10 pressure levels: $1000,925,850,700,500,300,200,100,50,10$ hpa, totalling 60 variables (full list in D.1.1).
ORAS5 or the Ocean Reanalysis System 5 provides an extensive record of sea-ice variables that incorporate multiple depth levels [29]. Since the public data is available on a monthly basis, we replicate them for daily compatibility with temporal extent from 1979 to present, for a total of 21 variables, including sst and ssh (full list in D.1.2).
LRA5 or ERA5-Land Reanalysis provides a detailed record of variables governing global terrestrial processes with specific corrections tailored for land surface applications such as flood forecasting [30] or carbon fluxes [21, 31]. We processed hourly data from 1979 to present and selected measurements at the 00UTC step, for a total of 43 variables, including t2m, u10, v10, and tp (full list in D.1.3).
@@ 3.3 Auxiliary
Figure 3: Motivating problem: as we perform longer rollouts, the (a) residual error becomes larger and prediction becomes blurry. This behavior is captured in the Fourier frequency domain where the (b) power spectra $S(k)$ at low wavenumber $k$ (i.e., low frequency signal) remains consistent at long rollouts, but not for higher $k$ (i.e., high frequency signal). This phenomenon explains why long-term forecasts excel at capturing large-scale pattern but not fine-grained details i.e., smooth.
@@ 4.2 Physics Metrics
As illustrated in Figure 3, we find that in general, data-driven forecasts tend to become blurry (Figure 3(a)) due to power divergence in the spectral domain (Figure 3(b) + S10). This motivates us to propose two physics-based metrics that measure the deviation or difference between the power spectra of prediction ${\hat{S}}(k)$ and target $S(k)$, where $k\in\mathbf{K}$, and $\mathbf{K}$ is the set of all scalar wavenumbers from 2D Fourier transform. Focusing on high-frequency components, we introduce $\mathbf{K}_{q}=\{k\in\mathbf{K}\mid k\geq Q(q)\}$, where $Q$ is the quantile function of $\mathbf{K}$ and $q\in[0,1]$. We set $q=0$ or $q=0.9$ for training and evaluation respectively. We denote $S_{q}=\{S(k)\mid k\in\mathbf{K}_{q}\}$ as the corresponding power spectra on $\mathbf{K}_{q}$, and we normalize the distribution to $S^{\prime}(k)$ such that it sums up to 1. Similarly we use $\hat{S}^{\prime}(k)$ to denote the normalized power for predictions.
(a) RMSE ($\downarrow$ is better)
(b) ACC ($\uparrow$ is better)
(c) MS-SSIM ($\uparrow$ is better)
Figure 4: Evaluation results between baseline climatology (black line) and physics-based control/deterministic forecasts. At longer forecasting horizon, most physics-based control/deterministic forecasts perform worse than climatology.
(a) RMSE ($\downarrow$ is better)
(b) ACC ($\uparrow$ is better)
(c) MS-SSIM ($\uparrow$ is better)
Figure 5: Evaluation results between baseline climatology (black line) and data-driven models including PanguWeather (PW), GraphCast (GC), and FourCastNetV2 (FCN2). We find that deterministic ML models perform worse than climatology on S2S timescale. Note: FCN2 lacks q-700.
@@ 5 Benchmark Results
Figure 7: Metrics ratio e.g., $\texttt{RMSE}_{ens}/\texttt{RMSE}_{det}$ between ensemble and deterministic forecasts, where the former improves the latter by accounting for IC uncertainty that can lead to trajectory divergences. Note: $n$ represents the number of ensemble members.
@@
Ensemble Forecasts Account for IC Uncertainty. Despite the underperformance of deterministic models, many studies have highlighted the potential of ensemble forecasts to account for trajectory divergences caused by IC uncertainties [44, 45, 46], also known as the butterfly effect [15]. Figure S4 shows that the performance of ensembles across physics-based models improves relative to their deterministic counterparts. For instance, when we take the metrics ratio between ensemble and deterministic forecasts as in Figure 7 (+ S5), the ratio of RMSE decreases with lead time, while the ratio of MS-SSIM improves over time with little significant changes in SpecDiv. The extent of improvement also appears to be affected by the number of ensemble members i.e., higher ensemble size $n$ appears to improve skillfulness. We also note similar insights from data-driven ensembling strategy as discussed in Section G.3. This highlights the importance of building a well-dispersed ensemble that accounts for long-range divergences for improved S2S predictability.
Figure 8: Probabilistic evaluation on ensemble forecasts indicating current skill limits of 15-20 days; CRPSS $>0$ suggests skills better than climatology variability. Note: $n$ represents the number of ensemble members.
Physical Constraints Yield Improved Performance. We find models that explicitly incorporate physical knowledge (e.g., learning spectral signals beyond pixel information) have better performance across metrics, such as FNO, as summarized in Table S8 given identical parameter budget of $10^{6}$. This phenomena is unsurprising and has been repeatedly demonstrated in many real-world applications of physics-informed deep learning, for instance.
Current Limits of S2S Predictability. Given our best models, we evaluate the extent of predictability in order to base our next steps. As illustrated in Figure 8 (+ S6), we find that ECMWF high-resolution ensemble, dubbed as the gold standard, still has the best performance in terms of CRPSS (vs ERA5 climatology), with a predictability range of around 15-20 days ahead before its skill collapses to climatology (i.e., CRPSS $\rightarrow$ 0). However, the resurgence of data-driven models are rapidly transforming the field as they are able to efficiently distil knowledge and automatically discover emergent patterns from large-scale, high-dimensional dataset, instead of first reducing them to physical functions with limited set of variables requiring constant calibration as is traditionally done in NWPs. The challenge, therefore, is to extend the predictability range of weather system as a representation of large-scale chaos, and we welcome the machine learning communities to take part in this open effort.
@@ 6 Conclusion
We present ChaosBench, a challenging benchmark to extend the predictability range of weather emulators into the S2S timescale where many processes with significant socioeconomic repercussions tend to occur, including extreme events. In addition to providing diverse datasets beyond ERA5 for a full Earth system emulation, we also perform extensive benchmarking on state-of-the-art data-driven and physics-based models alike. Through various ablation, we systematically find that skillfulness can be extended by ensemble forecasting, controlling for exponential error growth, and incorporating physical knowledge in our modeling approaches.
Future Work. Our input datasets have relatively coarse spatiotemporal resolution to match that of physics-based S2S forecasts. Nevertheless, we make the data processing pipeline open-source, allowing users to easily process inputs of the desired resolution (see Section B.4 for more details). We are planning for a multi-source reanalysis products (e.g., MERRA-2 [49]), leveraging diverse dataset strengths, such as the assimilation of different set of observations. As always, we welcome any contribution from the open-source community to solve this important yet understudied problem. And any comments, feedback, and/or future feature requests can be directed to the corresponding author or through the Github issue tracker at https://github.com/leap-stc/ChaosBench.
@@
We would like to thank Matthew Wilson, Tom Andersson, and Dale Durran for the insightful discussion during the earlier version of the manuscript. The authors also acknowledge funding, computing, and storage resources from the NSF Science and Technology Center (STC) Learning the Earth with Artificial Intelligence and Physics (LEAP) (Award #2019625) and the Department of Energy (DOE) Advanced Scientific Computing Research (ASCR) program (DE-SC0022255). AG would like to acknowledge support from Google and Schmidt Sciences. Last but definitely not least, we acknowledge the comprehensive S2S database emerging from the joint initiative of the World Weather Research Programme (WWRP) and the World Climate Research Programme (WCRP). The original S2S database is hosted at ECMWF as an extension of the TIGGE database.
@@
Juan Nathaniel1,*, Yongquan Qu1, Tung Nguyen2, Sungduk Yu3,5, Julius Busecke1,4,
@@
1Columbia University, 2UCLA, 3UCI, 4LDEO, 5 Intel Labs 11footnotetext: Corresponding author: jn2808@columbia.edu